OpenStudy (anonymous):

rewrite the equation x^2+y^2+3x=0 using the substitutions x=r cos thetha and y=r sin theta. simplify. please explain the steps!

OpenStudy (anonymous):

centre -3/2, 0 radius 3 so x+ 3/2 = 3cost y=3sint

OpenStudy (anonymous):

have you started out by plugging x=rcos(theta) and y=rsin(theta) into the equation? what does that give you?

OpenStudy (anonymous):

do you have to find the center and radius or is it done a different way?

OpenStudy (anonymous):

dont' worry about center. just plug x=rcos(theta) and y=rsin(theat) into x^2+y^2+3x=0. you'll get a cos^2 and a sin^2 that you can combine to get 1. try it and see what you get

OpenStudy (anonymous):

what do i do about the r?

OpenStudy (anonymous):

im sorry i haven't done this in a very long time

OpenStudy (anonymous):

\[(r \cos \theta)^{2} + (r \sin \theta)^{2} + 3r \cos \theta = 0\]\[r^{2}(\cos^{2}\theta + \sin^{2}\theta) + 3r \cos \theta = 0\]\[r^{2} + 3r \cos \theta = 0\]\[r(r + 3 \cos \theta) = 0\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

here...i'll get you started. :) abtrehearn got you most of the way.. realistically, you wouldn't have r=0 since the radius can't be 0 for a circle.

OpenStudy (anonymous):

so you could divide through by r.

OpenStudy (anonymous):

to get r+3cos\[\theta\]=0?

OpenStudy (anonymous):

sorry i messed up the format

OpenStudy (anonymous):

True. r = 0 is just a point. It does, however, lie on the graph of the original equation, \[r = -3 \cos \theta\]is the complete answer.